# Reflection Theory of Nichols Algebras over Coquasi-Hopf Algebras with Bijective Antipode

### v1

@@ 2.4 Yetter-Drinfeld modules in arbitrary braided monoidal category
By [Heckenberger and Schneider, [2020] ©2020, Proposition 3.4.8], $c^{YD}_{Y,X}$ is an isomorphism in $\mathcal{C}$ with inverse $\overline{c}_{X,Y}^{\mathcal{YD}}$.
The definition of the bosonization naturally extends to the case that $K$ is a Hopf algebra in the braided monoidal category $^{R}_{R}\mathcal{YD}(\mathcal{C})$}, where $\mathcal{C}$ is now an arbitrary braided monoidal category. We collect several key results from [Bespalov, 1997,Bespalov, 1995,Alonso Alvarez and Fernández Vilaboa, 2000] that will be important for the construction.
@@
3 Rational Yetter-Drinfeld module over coquasi-Hopf algebras
@@ 3.1 Rational modules in $^{H}_{H}\mathcal{YD}$}
Since $H$ is a non-associative algebra, we can not talk about modules over $H$ in the usual sense. However, if $R$ is an algebra in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$, then the category $_{R}\mathcal{C}$ of left $R$-modules in $\mathcal{C}$ is well-defined. Motivated by [Heckenberger and Schneider, [2020] ©2020, Definition 12.2.3], we now introduce the concept of rationality under the setting of coquasi-Hopf algebras.
@@
(1) A left module $(X,\lambda_{X})$ over $R$ in $\mathcal{C}$ is called rational if for any $x\in X$, there exists a natural number $n_{0}$ such that $\lambda(R(n)\otimes x)=0$ for all $n\geq n_{0}$. We denote the category of left rational module over $R$ in $\mathcal{C}$ by ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$.
(2) A left Yetter-Drinfeld module $X$ over $^{R}_{R}\mathcal{YD}(\mathcal{C})$} is called rational if $X$ is rational as a module of $R$ in $\mathcal{C}$. We denote the corresponding category by ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
@@
It is worth noting that for a locally finite $\mathbb{N}_{0}$-graded Hopf algebra $R$, this graded definition of rationality coincides with the classical one. Intuitively, the condition $\lambda(R(n)\otimes x)=0$ for large $n$ implies that the action on any element $x$ factors through a finite-dimensional truncation of $R$, which recovers the standard notion of rationality in the non-graded setting.
A natural question arises naturally, are rational modules closed under tensor products. The following lemma ensures this.
@@
Since $r^{2}_{-3}\in H$, and $V$ is an object in $^{H}_{H}\mathcal{YD}$}, we have $(r^{2}_{-3}\rhd v_{0})_{0}\in V$. The expression $\lambda_{V\otimes W}(r\otimes(v\otimes w))$ is a finite sum. Thus, by the definition of the rational module, there exists a natural number $n_{1}$ such that for $n\geq n_{1}$,
@@
This implies $V\otimes W\in{{}_{R}\mathcal{C}_{\operatorname{rat}}}$. Therefore, ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$ is a monoidal subcategory of $^{R}_{R}\mathcal{YD}(\mathcal{C})$}. The category ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$ is closed under arbitrary direct sums, subobjects and quotient objects.
@@ 3.2 Properties of pairings
Let $A$ and $B$ be objects in $\mathcal{C}$, and $\omega:A\otimes B\rightarrow\mathbbm{k}$ is a morphism in $\mathcal{C}$, named pairing. For subsets $X\subseteq A$ and $Y\subseteq B$, we define
@@
We say $\omega$ is non-degenerate, if $A^{\perp}=0$ and $B^{\perp}=0$.
@@
Let $E$ be a subobject of $A$, we need to show that $E^{\perp}$ is an object in $\mathcal{C}$. The proof of $E^{\perp}$ being a $H$-comodule is the same as in the Hopf algebra case, see [Heckenberger and Schneider, 2013a, Lemma 2.3].
To show $H\rhd E^{\perp}\subseteq E^{\perp}$. One need to verify that for any $b\in E^{\perp}$, $h\in H$ and $e\in E$,
@@
Here $b_{0}\in E^{\perp}$, because we have shown $E^{\perp}$ is a $H$-comodule. Also, $e_{0}^{i}\in E$ for all $i$. Note that $\omega$ is a morphism in $\mathcal{C}$, and the braiding $c$ is a natural isomorphism, we have
@@
We first show the map $G_{V}$ is well-defined and injective. Let $f\in\operatorname{Hom}_{C}(V,B\otimes W)$, and $g=G_{V}(f)$. For each $v\in V$, there is a finite-dimensional subobject $F\subseteq B$ with $f(v)=\sum_{i\in I}b^{i}\otimes w^{i}\in F\otimes W$. Now for any $a\in F^{\perp}$, by definition of $G_{V}$ and Lemma 3.4, we have
@@
By construction of $\langle V^{\prime}\rangle$, if we denote $V^{\prime}$ as $X_{0}$, this shows $g(F^{\perp}\otimes Y_{0})=0$. This property implies that for all $h\in H$, $v\in V^{\prime}$, we have $f(h\rhd v)\in F\otimes W$.
Recall that $X_{1}=\sum_{y\in Y_{0}}\operatorname{span}\{y_{0}\mid\delta_{V}(y)=y_{-1}\otimes y_{0}\},$ We are going to show $g(F^{\perp}\otimes X_{1})=0$. This is equivalent to verify $f(X_{1})\in F\otimes W$. Since $f$ is a morphism in $\mathcal{C}$, we have
$(\operatorname{id}\otimes f)\delta(h\rhd v)$ $=\delta(f(h\rhd v)).$
Recall $F$ is an object in $^{H}_{H}\mathcal{YD}$}, we know that $\delta(f(h\rhd v))\in\delta(F\otimes W)\subseteq H\otimes(F\otimes W)$. This implies $f((h\rhd v)_{0})\in F\otimes W.$ Hence $g(F^{\perp}\otimes X_{1})=0$. By iterated statements, we can prove that
$g(F^{\perp}\otimes\langle V\rangle)=0.$
@@
as an object in $\mathcal{C}$. Given $g\in\operatorname{Hom}_{\mathcal{C},\operatorname{rat}}(A\otimes V,W),$ let $g_{U}$ be the restriction of $g$ to $A\otimes U$, then there is a morphism $f_{U}:U\rightarrow B\otimes W$ with $G_{V}(f_{U})=g_{U}$. For $U_{1},U_{2}\in\mathcal{V}$, let $U_{1}\subseteq U_{2}$, we have $f_{U_{2}}\mid_{U_{1}}=f_{U_{1}}$, since $G_{V}$ is injective. Hence the maps $f_{U}$ deine a linear map $f:V\rightarrow B\otimes W$ by $f(v)=f_{U}(v)$, where $U$ is an element in $\mathcal{V}$ containing $v$. This construction yields $G_{V}(f)=g$. Thus $G_{V}$ is bijective for arbitrary $V$. ∎
@@ 3.3 Bijection between comodules and rational modules
We begin by introducing the notion of Hopf pairings. For our purpose, we assume that $A$ and $B$ are locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $\mathcal{C}$ in this section. Since this does not affect our study of Nichols algebras, we assume that $A$ and $B$ have bijective antipodes in $\mathcal{C}$.
@@
    If $\omega:A\otimes B\rightarrow\textbf{1}$ is a dual pair, then also
@@
    $\omega^{+\operatorname{cop}}:=\omega^{+}\circ(\operatorname{id}_{B}\otimes\mathcal{S}^{-1}_{A}):B^{\operatorname{cop}}\otimes A^{\operatorname{cop}}\rightarrow\mathbbm{k}.$ (3.11)
The following lemma holds in general boards of monoidal category, here we cite this lemma for simplicity.
@@
Let $(A,B,\omega)$ be a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algberas in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. The functor
@@
is a strict monoidal isomorphism of categories.
@@
Note that $\operatorname{Hom}_{C,\text{rat}}(A\otimes V,V)$ is the set of all $\lambda$ in $\operatorname{Hom}_{C}(A\otimes V,V)$ such that for all $v\in V$, there is a finite-dimensional object $F\subseteq B$, such that $\lambda(F^{\perp}\otimes v)=0.$ Since $B$ is locally finite, there is a natural number $n_{0}$ such that $F\subseteq\bigoplus_{i=0}^{n_{0}}B(i)$. Therefore $\lambda(A(n)\otimes v)=0$ for all $n>n_{0}$. Thus if $(V,\lambda)\in{{}_{A}\mathcal{C}}$, then $(V,\lambda)\in{{}_{A}\mathcal{C}_{\operatorname{rat}}}$ automatically.
@@
Now we write down explicit formula of $\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}$.
@@
For all $a,a^{\prime}\in A$, $v\in V$, we have $\lambda\circ(\operatorname{id}_{A}\otimes\lambda)(a\otimes(a^{\prime}\otimes v))=\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}(a\otimes(a^{\prime}\otimes v)).$ By non-degeneracy of $\omega$ and Mac-Lane’s coherence theorem, this is equivalent to
@@ 3.4 Monoidal equivalence between Yetter-Drinfeld module categories related by a dual pair
Let $(A,B,\omega)$ be a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algberas with bijective antipodes in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. The main result in this subsection is to show there is a braided monoidal equivalence $(\Omega,\beta):{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
@@
where $D_{1},D_{2}$ is given by Lemma 3.9, then we have
@@
Here the second equality uses (3.9) and (3.11), the third and the sixth equalities use naturality of braiding, the fourth and fifth equalities use braiding axiom. ∎
Now we return to the setting of coquasi-Hopf algebras. We want to restrict the Yetter-Drinfeld module criterion to rational modules. Let $B$ be a locally finite $\mathbb{N}_{0}$-graded Hopf algbera in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ with bijective antipode. Let $(V,\lambda)\in{{}_{B}\mathcal{C}_{\operatorname{rat}}}$, and $(V,\delta)\in{{}^{B}\mathcal{C}}$.
@@
$S_{2}=\{(\overline{\lambda},\overline{\delta})\mid(V,\overline{\lambda})\in{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}},(V,\overline{\delta})\in{{}^{A^{\operatorname{cop}}}\overline{\mathcal{C}}}\}$
@@
$\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$${{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$$\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$$\overline{{\mathcal{YD}(\mathcal{C})^{A}_{A}}_{,\operatorname{rat}}}$$\scriptstyle{\Gamma}\scriptstyle{\cong}$$\scriptstyle{\Omega}$$\scriptstyle{F_{lr}}$$\scriptstyle{(F_{rl},\alpha)}$
@@
s depicted in the diagram above, we define the functor $\Omega$ as the composition:
@@
We now compute the explicit structure of this compositionLet $(V,\lambda_{B},\delta_{B})\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$. Then
@@
Thus $(\Omega,\overline{\beta}):\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}\rightarrow\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$ is a braided monoidal equivalence. Now we take reverse braiding, and it is direct that

### v2

@@ 2.4 Yetter-Drinfeld modules in arbitrary braided monoidal categories
By [Heckenberger and Schneider, [2020] ©2020, Proposition 3.4.8], $c^{\mathcal{YD}}_{Y,X}$ is an isomorphism in $\mathcal{C}$ with inverse $\overline{c}_{X,Y}^{\mathcal{YD}}$.
The definition of the bosonization naturally extends to the case where $K$ is a Hopf algebra in the braided monoidal category $^{R}_{R}\mathcal{YD}(\mathcal{C})$}, where $\mathcal{C}$ is now an arbitrary braided monoidal category. We collect several key results from [Bespalov, 1997,Bespalov, 1995,Alonso Alvarez and Fernández Vilaboa, 2000] that will be important for the construction.
@@
3 Rational Yetter-Drinfeld modules over coquasi-Hopf algebras
@@ 3.1 Rational modules in $^{H}_{H}\mathcal{YD}$}
Since $H$ is a non-associative algebra, we cannot talk about modules over $H$ in the usual sense. However, if $R$ is an algebra in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$, then the category $_{R}\mathcal{C}$ of left $R$-modules in $\mathcal{C}$ is well-defined. Motivated by [Heckenberger and Schneider, [2020] ©2020, Definition 12.2.3], we now introduce the concept of rationality in the setting of coquasi-Hopf algebras.
@@
(1) A left module $(X,\lambda_{X})$ over $R$ in $\mathcal{C}$ is called rational if for any $x\in X$, there exists a natural number $n_{0}$ such that $\lambda(R(n)\otimes x)=0$ for all $n\geq n_{0}$. We denote the category of left rational modules over $R$ in $\mathcal{C}$ by ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$.
(2) A left Yetter-Drinfeld module $X$ in $^{R}_{R}\mathcal{YD}(\mathcal{C})$} is called rational if $X$ is rational as a module of $R$ in $\mathcal{C}$. We denote the corresponding category by ${{}^{R}_{R}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
@@
We say that a $\mathbb{N}_{0}$-graded Hopf algebra $R$ is locally finite if $R(n)$ is finite-dimensional for all $n\geq 0$. It is worth noting that for a locally finite $\mathbb{N}_{0}$-graded Hopf algebra $R$, this graded definition of rationality coincides with the classical one. Intuitively, the condition $\lambda(R(n)\otimes x)=0$ for large $n$ implies that the action on any element $x$ is nontrivial on a finite-dimensional truncation of $R$, which recovers the standard notion of rationality in the non-graded setting.
A natural question is whether rational modules are closed under tensor products. The following lemma ensures this.
@@
Since $r^{2}_{-3}\in H$, and $V$ is an object in $^{H}_{H}\mathcal{YD}$}, we have $(r^{2}_{-3}\rhd v_{0})_{0}\in V$.
The expression $\lambda_{V\otimes W}(r\otimes(v\otimes w))$ is a finite sum. Thus, by the definition of the rational module, there exists a natural number $n_{1}$ such that for $n\geq n_{1}$,
@@
This implies $V\otimes W\in{{}_{R}\mathcal{C}_{\operatorname{rat}}}$. Therefore, ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$ is a monoidal subcategory of $^{R}_{R}\mathcal{YD}(\mathcal{C})$}. The category ${{}_{R}\mathcal{C}}_{\operatorname{rat}}$ is obviously closed under arbitrary direct sums, subobjects and quotient objects.
@@ 3.2 Properties of pairings
Let $A$ and $B$ be objects in $\mathcal{C}$, and $\omega:A\otimes B\rightarrow\mathbbm{k}$ is a morphism in $\mathcal{C}$, called a pairing. For subsets $X\subseteq A$ and $Y\subseteq B$, we define
@@
We say $\omega$ is non-degenerate if $A^{\perp}=0$ and $B^{\perp}=0$.
@@
Let $E$ be a subobject of $A$, we need to show that $E^{\perp}$ is an object in $\mathcal{C}$. The proof of $E^{\perp}$ being an $H$-comodule is the same as in the Hopf algebra case, see [Heckenberger and Schneider, 2013a, Lemma 2.3].
To show $H\rhd E^{\perp}\subseteq E^{\perp}$. One needs to verify that for any $b\in E^{\perp}$, $h\in H$ and $e\in E$,
@@
Here $b_{0}\in E^{\perp}$, because we have shown $E^{\perp}$ is an $H$-comodule. Also, $e_{0}^{i}\in E$ for all $i$. Note that $\omega$ is a morphism in $\mathcal{C}$, and the braiding $c$ is a natural isomorphism, we have
@@
We first show the map $G_{V}$ is well-defined and injective. Let $f\in\operatorname{Hom}_{C}(V,B\otimes W)$, and $g=G_{V}(f)$. By assumption, for each $v\in V$, there is a finite-dimensional subobject $F\subseteq B$ with $f(v)=\sum_{i\in I}b^{i}\otimes w^{i}\in F\otimes W$. Now for any $a\in F^{\perp}$, by definition of $G_{V}$ and Lemma 3.4, we have
@@
By construction of $\langle V^{\prime}\rangle$ in Lemma 3.5, if we denote $V^{\prime}$ as $X_{0}$, this shows $g(F^{\perp}\otimes Y_{0})=0$.
Recall that $X_{1}=\sum_{y\in Y_{0}}\operatorname{span}\{y_{0}\mid\delta_{V}(y)=y_{-1}\otimes y_{0}\},$ We are going to show $g(F^{\perp}\otimes X_{1})=0$. Since $g$ is a morphism in $\mathcal{C}$, we have
$(\operatorname{id}_{H}\otimes g)\circ\delta(F^{\perp}\otimes Y_{0})=\delta(g(F^{\perp}\otimes Y_{0})))=0.$
By definition of $X_{1}$,the set $\operatorname{span}\{z_{0}\mid\delta(z)=z_{-1}\otimes z_{0},\ z\in F^{\perp}\otimes Y_{0}\}=F^{\perp}\otimes X_{1}$. This implies $g(F^{\perp}\otimes X_{1})=0$. By iterating this argument, we can prove that
$g(F^{\perp}\otimes\langle V^{\prime}\rangle)=0.$
@@
as an object in $\mathcal{C}$. Given $g\in\operatorname{Hom}_{\mathcal{C},\operatorname{rat}}(A\otimes V,W),$ let $g_{U}$ be the restriction of $g$ to $A\otimes U$, then there is a morphism $f_{U}:U\rightarrow B\otimes W$ with $G_{V}(f_{U})=g_{U}$. For $U_{1},U_{2}\in\mathcal{V}$, let $U_{1}\subseteq U_{2}$, we have $f_{U_{2}}\mid_{U_{1}}=f_{U_{1}}$, since $G_{V}$ is injective. Hence the maps $f_{U}$ define a linear map $f:V\rightarrow B\otimes W$ by $f(v)=f_{U}(v)$, where $U$ is an element in $\mathcal{V}$ containing $v$. This construction yields $G_{V}(f)=g$. Thus $G_{V}$ is bijective for arbitrary $V$. ∎
@@ 3.3 Bijection between comodules and rational modules
We begin by introducing the notion of Hopf pairings. For our purposes, we assume that $A$ and $B$ are locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ in this section. Since this does not affect our study of Nichols algebras, we assume that $A$ and $B$ have bijective antipodes in $\mathcal{C}$.
@@
    If $\omega:A\otimes B\rightarrow\textbf{1}$ is a dual pair, then
@@
    $\omega^{+\operatorname{cop}}:=\omega^{+}\circ(\operatorname{id}_{B}\otimes\mathcal{S}^{-1}_{A}):B^{\operatorname{cop}}\otimes A^{\operatorname{cop}}\rightarrow\mathbbm{k}$ (3.11)
    are dual pairs too.
Recall that $\overline{C}$ means the reverse braided monoidal category. The following lemma holds in general braided monoidal categories; here we cite this lemma for simplicity.
@@
Let $(A,B,\omega)$ be a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algebras in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. The functor
@@
is a strict monoidal isomorphism.
@@
Note that $\operatorname{Hom}_{C,\text{rat}}(A\otimes V,V)$ is the set of all $\lambda$ in $\operatorname{Hom}_{C}(A\otimes V,V)$ such that for all $v\in V$, there is a finite-dimensional object $F\subseteq B$, such that $\lambda(F^{\perp}\otimes v)=0.$ Since $B$ is locally finite, there is a natural number $n_{0}$ such that $F\subseteq\bigoplus_{i=0}^{n_{0}}B(i)$. Therefore $\lambda(A(n)\otimes v)=0$ for all $n>n_{0}$. Thus if $(V,\lambda)\in{{}_{A}\mathcal{C}}$, where $\lambda\in\operatorname{Hom}_{C,\text{rat}}(A\otimes V,V)$, then $(V,\lambda)\in{{}_{A}\mathcal{C}_{\operatorname{rat}}}$ automatically.
@@
Now we write down explicit formula of $\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}$
@@
Since we have assumed $(V,\lambda)\in{{}_{A}\mathcal{C}}$, for all $a,a^{\prime}\in A$, $v\in V$, we have $\lambda\circ(\operatorname{id}_{A}\otimes\lambda)(a\otimes(a^{\prime}\otimes v))=\lambda\circ(\mu_{A}\otimes\operatorname{id}_{V})\circ a_{A,A,V}^{-1}(a\otimes(a^{\prime}\otimes v)).$ By non-degeneracy of $\omega$ and Mac-Lane’s coherence theorem, this is equivalent to
@@ 3.4 Monoidal equivalence between Yetter-Drinfeld module categories related by a dual pair
Let $(A,B,\omega)$ be a dual pair of locally finite $\mathbb{N}_{0}$-graded Hopf algebras with bijective antipodes in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. The main result in this subsection is to show there is a braided monoidal equivalence $(\Omega,\beta):{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$.
@@
where $D_{1},D_{2}$ is given by Lemma 3.9, then
@@
Here the second equality uses (3.9) and (3.11), the third and the sixth equalities use the naturality of the braiding, the fourth and the fifth equalities use the braiding axiom. ∎
Now we return to the setting of coquasi-Hopf algebras. We want to restrict the Yetter-Drinfeld module criterion to rational modules. Let $B$ be a locally finite $\mathbb{N}_{0}$-graded Hopf algebra in $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$ with bijective antipode. Let $(V,\lambda)\in{{}_{B}\mathcal{C}_{\operatorname{rat}}}$, and $(V,\delta)\in{{}^{B}\mathcal{C}}$.
@@
$S_{2}=\{(\overline{\lambda},\overline{\delta})\mid(V,\overline{\lambda})\in{{}_{A^{\operatorname{cop}}}\overline{\mathcal{C}}_{\operatorname{rat}}},(V,\overline{\delta})\in{{}^{A^{\operatorname{cop}}}\overline{\mathcal{C}}}\}.$
@@
$\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$${{}^{A^{\operatorname{cop}}}_{A^{\operatorname{cop}}}\mathcal{YD}(\overline{\mathcal{C}})}_{\operatorname{rat}}$$\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$$\overline{{\mathcal{YD}(\mathcal{C})^{A}_{A}}_{,\operatorname{rat}}}$$\scriptstyle{\Gamma}\scriptstyle{\cong}$$\scriptstyle{\Omega}$$\scriptstyle{F_{lr}}$$\scriptstyle{\overline{F_{rl}}}$
@@
As depicted in the diagram above, we define the functor $\Omega$ as the composition:
@@
We now compute the explicit structure of this composition. Let $(V,\lambda_{B},\delta_{B})\in{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}$. Then
@@
Thus $(\Omega,\overline{\beta}):\overline{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}\rightarrow\overline{{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}}$ is a braided monoidal equivalence. Now we take the reverse braiding, and it is direct that
$(\Omega,\beta):{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}}\rightarrow{{}^{A}_{A}\mathcal{YD}(\mathcal{C})}_{\operatorname{rat}},\ \beta_{X,Y}=c_{Y,X}^{{{}^{B}_{B}\mathcal{YD}(\mathcal{C})}}\circ\overline{c}_{X,Y}$